OpenStudy (anonymous):

Greatest common factor of 4, 16

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4

OpenStudy (aroub):

4

OpenStudy (anonymous):

4=4*1 16=4*4 GCF=4

OpenStudy (anonymous):

The GCF of 4 and 16 is 4. 4 = 1 x 4 16 = 4 x 4

OpenStudy (anonymous):

thank you. my baby brother will appreciate it :)